[Hallb] [Revised Logentry] RC Summary from (7 PM Oct. 3rd to 8 AM Oct. 4th)

elfassi at jlab.org elfassi at jlab.org
Wed Oct 4 20:20:01 EDT 2023

Logentry Text:
Oct. 3rd. Day - Swing Shifts:

After the target was signed off around 1 PM, the run certificate was issued so MCC could get Hall B ready for beam tuning. It took MCC a couple of hours to power-cycle our beamline power supplies and set up the beam orbits. 

6 - 7 PM: MCC starts sending a pulse beam.

7 PM: Got CW beam to the tagger dump and requested 5 nA for harp scans; 2C21A followed by 2C24A. See their results, respectively, at https://logbooks.jlab.org/entry/4191945 and https://logbooks.jlab.org/entry/4191955. 

8 - 9 PM: Several hiccups happened starting from the solenoid trip due to the LCW issue and low supply pressure; see the logbook entry #: https://logbooks.jlab.org/entry/4192030, IOCCLASSC1 crashed twice, and the Moller DAQ once we configure and start a Moller run. 

9 PM - 12 AM: Dedicated to understanding the failure of the Moller run due to various reasons, such as the splitter that was added to get ADC rates even after accessing the hall to remove it, high rates of accidentals compared to coincidences, etc. We also waited for the Magnet group to debug the solenoid issue reported to them, which led to the conclusion that a prompt controlled access is needed to debug it. Carroll Jones and our engineering-n-call Morgan were called in to resolve this issue and avoid losing much beam time. 

Oct. 4th Owl Shift: 

12 - 3 AM: Started with the LCW/Solenoid controlled access is about to take place, which resulted in the issue resolution, as reported on https://logbooks.jlab.org/entry/4192030. Based on Probir's recommendations, the solenoid current was gradually ramped up by starting with 100 A, watching the LCW behavior for 1/2h, ramping up to 250 A, and watching for another 1/2 before switching it to the 1/2 solenoid field intensity. Eugene was called to assist with that. Meanwhile, we asked MCC to degauss the tagger magnet and mask the beam halos; then, we moved the collimator to 20 mm.

3 - 5 AM: Once the current corresponding to 1/2 field-intensity solenoid was reached, the process of tuning the beam to the Faraday cup started by performing the 2H01 harp scan; see the nice shift summary of Izzy for detailed performed tasks, https://logbooks.jlab.org/entry/4192064, as well as https://logbooks.jlab.org/entry/4192230. 

5 - 8 AM: Spent on centering beam on target using the empty cell exit and entrance windows to determine new BPMs X & Y positions, especially for 2H01, as well as the new FSD thresholds. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4192458
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallb/attachments/20231004/2f68e796/attachment.html>

More information about the Hallb mailing list